Reports 1-1 of 1 Clear search Modify search
AdV-PSL (Commissioning of SS)
derossi, gosselin - 15:08 Thursday 19 February 2026 (68717) Print this report
ML cristal temperature investigation

In order to find a good crystal temperature setpoint, working both for the master laser and the squeezing laser, we investigated temperatures from the current value, which is 23.7C, up to 31.4C.

The mode hops appears as an additional error signal while scanning the SL piezo.

ML crystal temp gps UTC considerations  
26.8 C  10:06:00    
26.4 C  10:08:00    
26.0 C  10:09:00    
25.7 C 10:11:00 we observed a peak appearing close to the error signal which, however is not a mode hop  
25.8 C 10:12:30    
25.5 C  10:16:00 the error signal starts to worsen  
25.3 C 10:20:00 a second error signal appeared (mode hop, probably the same mentioned in https://logbook.virgo-gw.eu/virgo/?r=67348 fig. 1
24.8 C   10:23:00 the mode hop and the fundamental mode error signals have the same amplitude fig.2
24.6 C 10:24:00 the mode hop disappear   
23.7 C   10:28:00  this is the nominal temperature. However some peaks equally spaced are visible (noise at the 14MHz harmonics?) (fig.3 and 6)  
27.3 C 10:33: 00 it seems a good setpoint  fig.4
27.9 C 10:35:00 faint error signal  
28.1 C 10:37:00  mode hop  fig. 5
29.0 C 10:39:00 another good setpoint  
31.4 C 10:44:00 mode hop   

 

In conclusion, a temperature around 29.7 C seems the best candidate since there are no spurious peaks and it is in the middle betwwen two mode hops. Also 27.3 C has a good shape but it is closer to the 28.1 C mode hop.
 

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×